OpenStudy (anonymous):

\(x^3 -2x^2 + a \equiv \left(x+2\right) Q(x) + 3 \) where Q(x) is a polynomial. Find the value of a

OpenStudy (anonymous):

Well, how do i start this? i havent done these in a while..

OpenStudy (anonymous):

if \(x=-2\) the right hand side is \(3\) make sure the left hand side is 3 as well

OpenStudy (anonymous):

\[(-2)^2 -2(-2)^2+a \equiv \left(x+2\right) Q(x) + 3 \] hmm \[-16 + a \equiv \left(x+2\right) Q(x) + 3 \]

OpenStudy (anonymous):

right hand side is 3

OpenStudy (anonymous):

how is the right hand side 3? like this\[(-3+2)(Q(-3) +3\]?

OpenStudy (anonymous):

\[(-2)^2 -2(-2)^2+a \equiv \left(-2+2\right) Q(-2) + 3\]

OpenStudy (anonymous):

ohhh..i got it thanks!

OpenStudy (anonymous):

yw